OpenStudy (anonymous):

Subtract 9/2 - 7/-5 Write answer in simplest form

OpenStudy (debbieg):

So you might want to start by dealing with the negative in the 2nd denominator, just to make things easier to look at. Pull it out in front and what does that do to the operator (subtraction) in the middle of the terms? After that, you need an LCD. What will that be?

OpenStudy (anonymous):

14

OpenStudy (debbieg):

The LCD? No, not 14... how did you get that?

OpenStudy (anonymous):

2nd fraction gets turned around does it not?

OpenStudy (debbieg):

\[\frac{ 9 }{ 2 }-\frac{ 7 }{ -5 }=\frac{ 9 }{ 2 }+\frac{ 7 }{ 5 }\] now what? "Turned around" as in flip it, take the reciprocal? No... you may be thinking of division?

OpenStudy (anonymous):

Oh my bad so the LCD would be 10

OpenStudy (debbieg):

OK, good.... now you have an LCD. What does that make your numerators? rewrite the problem with the LCD of 10.

OpenStudy (anonymous):

\[\frac{ 45 }{ 10 } \frac{ -14 }{ 10 }= \frac{ 59 }{ 10 }\]

OpenStudy (debbieg):

Very good. :)

OpenStudy (anonymous):

What about \[\frac{ 8 }{ -9 }+\frac{ 7 }{ 6 }\]

OpenStudy (debbieg):

Again, you need LCD. The LCD this time is a bit different - what do you think it is?

OpenStudy (debbieg):

(and if that negative denominator annoys you, you could pull it out in front, or even swap the order of the terms and change it to subtraction, e.g., -A+B=B-A)

OpenStudy (anonymous):

18

OpenStudy (debbieg):

Good. so again, build up each numerator over that LCD.

OpenStudy (anonymous):

Ok, but what if i have a negative on the denominator of the first fraction and a negative numerator on the second?

OpenStudy (debbieg):

Well, the sign of one term doesn't have anything to do with the other. In that case, both of the terms of negative. If the operator is subtraction, that would cancel with the negative in the second term, though. E.g, if a, b, c, and d are all positive: \[\frac{ a }{-b }+\frac{ -c }{d }=-\frac{ a }{b }-\frac{ c }{d }\]\[\frac{ a }{-b }-\frac{ -c }{d }=-\frac{ a }{b }+\frac{ c }{d }\]

OpenStudy (anonymous):

oh ok thank you

OpenStudy (debbieg):

And don't forget: -a-b=-(a+b) and -a+b=b-a You can shuffle/rearrange the terms to make it easier to evaluate, as long as you obey the rules!